OpenStudy (steve816):

Please help on some trigonometry, it would be greatly appreciated! How to solve this equation?

OpenStudy (steve816):

\[\sin (3\theta) = -1\]

OpenStudy (baru):

can you slove this? sin x = -1

OpenStudy (steve816):

yeah, that would be 3pi/2

OpenStudy (baru):

now make the substitution \(3\theta =x \)

OpenStudy (baru):

\(\theta = ?\)

OpenStudy (steve816):

3pi/6

OpenStudy (baru):

yes, look at the fraction 3pi/6 you can simplify it further

OpenStudy (steve816):

pi/2

OpenStudy (baru):

yep

OpenStudy (owlcoffee):

What we can learn from this is pretty much that whenever solving a trigonometric equation and the given form does not fit in any of the trigonometric entities or equalities then we MUST make a change of variable in order to make it fit and later on solve for the original variable.